A wire is stretched from the top of a 12 foot pole to a stake 5 feet from the base of the pole. How long was the wire?

Answer by checkley77(12844)

A^2+B^2=C^2
12^2+5^2=C^2
144+25=C^2
169=C^2
C=SQRT169
C=13 FEET ANS,